OpenStudy (anonymous):

56x^2-113x+56 factors and show the steps how u got the answer please

OpenStudy (anonymous):

this does not factor jalesia - u sure u got the correct numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the middle number 112?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

7x -8)(8x -7)

OpenStudy (anonymous):

There

OpenStudy (anonymous):

right - sorry - i missed that!!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

The steps are to simply look for factors that multiply to get 56, but add up to get 113...you just keep trying until you get it

OpenStudy (anonymous):

\[\sqrt{64x^2y^2}\]

OpenStudy (anonymous):

Take the sqrt of each number and variable to get 8xy

OpenStudy (anonymous):

show the step how u got the answer

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[\sqrt{64x^2y^2} = \sqrt{64} \sqrt{x^2} \sqrt{y^2}\]

OpenStudy (anonymous):

sqrt 64 = 8 sqrt x^2 = x sqrt y^2 = y

OpenStudy (anonymous):

And so the answer is 8xy

OpenStudy (anonymous):

When you're doing these you have to break them into parts - first take the square root of the number, then the first variable, then the second variable

OpenStudy (anonymous):

jalesia, are you still here?

OpenStudy (anonymous):

Do you have any questions?